Problem 3-11 (Algo) T-Account Analysis of Cost Flows [LO3-2, LO3-3, LO3-4) Selected T-accounts of Moore Company are given belComplete this question by entering your answers in the tabs below. Req 1 to 5 Req 6 Req 7 Req 8 1. What was the cost of raw mComplete this question by entering your answers in the tabs below. Req 1 to 5 Reg 6 Reg 7 Reg 8 If overhead is applied to pro$10,750 of this balance is direct labor cost, how much of it is direct materials cost? Applied overhead cost? Complete this qComplete this question by entering your answers in the tabs below. Req 1 to 5 Reg 6 Reg 7 Reg 8 Compute the ending balance in
Answer #1

1) Cost of raw material used

= Beginning +purchased-ending

= 26000 + 142000 - 36000

= $132000

2) Indirect material

= total material -indirect material

= 132000 - 101000

= $ 31000..

3) Indirect labor

= Current factory cost-direct labor

= 202000 - 183000

= $ 19000

4) Cost of goods manufactured : $503000 (credited in work in process)

5) Unadjusted Cost of goods sold

= Beginning FG+cost of good manufactured-ending FG

= 51000 + 503000 - 76500

= $ 477500

6) Overhead rate

= Overhead /direct labor cost

= 223260 / 183000

=1.22 or 122% of direct labor cost

7) Manufacturing overhead is overapplied as actual overhead is less than applied : 223260 - 202260 = 21000

8) Ending WIP : 31000 + 101000 + 183000 + 223260 - 503000 = 35260

Applied overhead : 10750*122% = $ 13115

Direct material

= Balance - labor - overhead

= 35260 - 10750 - 13115

= $ 11395